# Inner crust of neutron stars: Unified description of structure and superconductivity

### v1

@@ C.2 Variational equations
where $y$ is either of the variables from the set $\{n,\,x,\,n_{no},\,r_{B},\,u^{\rm bub},\,n_{sk}^{\rm bub},\,n_{dc}^{\rm bub}\}$.
@@ C.3 Optimum size of bubble
The derivative
$\left.\frac{\partial(w_{\rm tot}^{\rm bub})}{\partial r_{B}}\right|_{n,x,n_{no},u^{\rm bub},n_{sk}^{\rm bub},n_{dc}^{\rm bub}}=0$ (151)
@@
Notice a difference in signs in Eqs. (79) and (153). The solution to Eq. (153) can be readily found using the methods of Ferrari and Cardano resulting in
@@ C.4 Chemical equilibrium
The derivative
$\left.\frac{\partial(w_{\rm tot}^{\rm bub})}{\partial x}\right|_{n,n_{no},r_{B},u^{\rm bub},n_{sk}^{\rm bub},n_{dc}^{\rm bub}}=0$ (159)
@@ C.5 Continuity of neutron chemical potential across the bubble interface
The derivative
$\left.\frac{\partial(w_{\rm tot}^{\mathrm{bub}})}{\partial n_{ni}}\right|_{N_{n},n_{pi},r_{B},u^{\mathrm{bub}},n_{sk}^{\mathrm{bub}},n_{dc}^{\mathrm{bub}}}=0$ (163)
@@
The partial derivative is given by:
$\left.\frac{\partial}{\partial n_{ni}}\right|_{N_{n},\,n_{pi},\,r_{B},\,u^{\mathrm{bub}},n_{sk}^{\mathrm{bub}},n_{dc}^{\mathrm{bub}}}=$ (166)
$\left.\frac{\partial}{\partial n}\right|_{x,\,n_{no},\,r_{B},\,u^{\mathrm{bub}},n_{sk}^{\mathrm{bub}},n_{dc}^{\mathrm{bub}}}+\frac{\delta x}{\delta n}\left.\frac{\partial}{\partial x}\right|_{n,\,n_{no},\,r_{B},\,u^{\mathrm{bub}},n_{sk}^{\mathrm{bub}},n_{dc}^{\mathrm{bub}}}$
$+\frac{\delta n_{no}}{\delta n_{ni}}\left.\frac{\partial}{\partial n_{no}}\right|_{n,\,x,\,r_{B},\,u^{\mathrm{bub}},n_{sk}^{\mathrm{bub}},n_{dc}^{\mathrm{bub}}}.$
@@
$\left.\frac{\partial}{\partial n_{ni}}\right|_{N_{n},\,n_{pi},\,r_{B},\,u^{\mathrm{bub}},n_{sk}^{\mathrm{bub}},n_{dc}^{\mathrm{bub}}}=$ (167)
$\left.\frac{\partial}{\partial n}\right|_{x,\,n_{no},\,r_{B},\,u^{\mathrm{bub}},n_{sk}^{\mathrm{bub}},n_{dc}^{\mathrm{bub}}}-\frac{x}{n}\left.\frac{\partial}{\partial x}\right|_{n,\,n_{no},\,r_{B},\,u^{\mathrm{bub}},n_{sk}^{\mathrm{bub}},n_{dc}^{\mathrm{bub}}}$
$-\frac{1-u^{\mathrm{bub}}}{u^{\mathrm{bub}}}\left.\frac{\partial}{\partial n_{no}}\right|_{n,\,x,\,r_{B},\,u^{\mathrm{bub}},n_{sk}^{\mathrm{bub}},n_{dc}^{\mathrm{bub}}}.$
@@
$+\frac{1-x}{(1-u^{\mathrm{bub}})n}\left.\frac{\partial(w_{\rm p.surf}^{\rm bub}+w_{\rm curv}^{\rm bub})}{\partial x}\right|_{r_{B},u^{\mathrm{bub}},n_{sk}^{\mathrm{bub}},n_{dc}^{\mathrm{bub}}}$
$+\frac{1}{(1-u^{\mathrm{bub}})n}\left.\frac{\partial(w_{\rm C+L}^{\rm bub})}{\partial x}\right|_{n,r_{B},u,n_{sk}^{\mathrm{bub}},n_{dc}^{\mathrm{bub}}}.$ (171)
@@ C.6 Continuity of pressure across the bubble interface
The derivative
$\left.\frac{\partial(w_{\rm tot}^{\mathrm{bub}})}{\partial u^{\mathrm{bub}}}\right|_{N_{n},N_{no},x,V_{c},n_{sk}^{\mathrm{bub}},n_{dc}^{\mathrm{bub}}}=0$ (172)
implies that the energy density is minimized by choosing the optimal size of the nucleus while keeping fixed the cell size, the number of neutrons inside the nucleus, the number of neutrons outside the nucleus and the proton fraction. As a result, the pressure in the nucleus is equal to the pressure in the dripped neutrons, $P_{i}^{\mathrm{bub}}=P_{o}^{\mathrm{bub}}$.
@@
The partial derivative is given by:
@@
×$\times(w_{\rm p.surf}^{\mathrm{bub}}+w_{\rm cur}^{\mathrm{bub}}+w_{\rm C+L}^{\mathrm{bub}}),$
@@
×$\times\left[\left(\frac{2+d}{d}+\frac{2u^{\mathrm{bub}}}{1-u^{\mathrm{bub}}}\right)f_{d}(u^{\mathrm{bub}})+u^{\mathrm{bub}}\partial_{u^{\mathrm{bub}}}f_{d}(u^{\mathrm{bub}})\right].$
C.7 Optimal energy to add neutron to the skin or to the density curvature region
Analogously to the case of nuclei, the conditions
$\left.\frac{\partial(E_{\rm tot}^{\rm bub}/V_{c})}{\partial n_{sk}^{\rm bub}}\right|_{n,x,N_{n},r_{B},u^{\rm bub},n_{dc}^{\rm bub}}=0$ (187)
and
$\left.\frac{\partial(E_{\rm tot}^{\rm bub}/V_{c})}{\partial n_{dc}^{\rm bub}}\right|_{n,x,N_{n},r_{B},u^{\rm bub},n_{sk}^{\rm bub}}=0$ (188)
simply lead to
$\mu_{sk}=\mu_{no}^{\rm bub},$ (189)
$\mu_{dc}=\mu_{no}^{\rm bub}.$ (190)
@@ Appendix D Nuclear surface tension with curvature corrections
To obtain self-consistently the nuclear surface tension with curvature corrections, let us start considering a neutron-rich nucleus in equilibrium with a liquid of dripped neutrons. The problem of finding the density profile is formally similar to the one found in hot nucleus in equilibrium with the pure liquid phase. It is known that in this scenario the variational equations at given temperature and chemical potential possess two different solutions in both, quantum bonche85 and semiclassical descriptions suraud87 . One of these solutions represents the nucleus-plus-liquid system and the another one to the pure liquid. Therefore, it seems natural to define the extensive quantities in the nucleus as the differences of their values in the nucleus-plus-liquid and the pure-liquid system.
To simplify this calculation, we simulate the nucleus as a half-space filled by uniform nuclear matter ($\tilde{z}<0$), while the other half-space is filled by the pure neutron liquid ($\tilde{z}>0$). As a consequence, in this geometry the energy per unit area, i.e. the surface tension, of nucleus constrained by the conservation of the neutron and protons numbers reads Vinas1998 :
[…]
where $\varepsilon_{d}$ is the energy density of the neutron liquid and $\varepsilon(\tilde{z})$ is the local energy density of the nucleus plus pure liquid (LG) system, which in the $\tilde{z}\to-\infty$ corresponds to the energy density $\varepsilon_{0}$ of uniform asymmetric nuclear matter in equilibrium with the uniform liquid of dripped neutrons (G). Here, $n_{n}(\tilde{z})$ is the total number density of neutrons at the plane $\tilde{z}=\mathrm{const}$, $n_{no}$ is the number density of neutrons of the pure neutron liquid at $\tilde{z}\to+\infty$ and $n_{p}=n_{p}(\tilde{z})$ is the number density of the proton component of nuclear matter.
The neutron and proton density profiles as well as the density of dripped neutrons are the solution of the Euler-Lagrange equations associated to Eq.(191):
where $\mu_{n}$ and $\mu_{p}$ are the neutron and proton chemical potentials computed at $\tilde{z}\to-\infty$, i.e. corresponding to the bulk phase in equilibrium with the dripped neutron liquid, which due to the equilibrium has the same neutron chemical potential, and $\varepsilon_{no}=\varepsilon(n=n_{no},x=0)$. The Euler-Lagrange equations (192) is a set of coupled second order differential equations that are solved using the imaginary time-step method and provide the self-consistent neutron and proton densities in the LG and G phases. It is worthwhile to mention here that the surface tension defined by Eq. (191) corresponds to the surface contribution to the thermodynamical potential $\Omega=E-\mu_{n}N-\mu_{p}Z=-PV$ in the LG phase, which is the quantity that has to be minimum in the thermodynamical equilibrium between the LG and G phases douchin00 .
The surface tension in Eq. (191) has been obtained under the assumption of a planar interface. In order to take into account that the interface is curved, we add an additional curvature correction that takes into account the curvature of the interface douchin00 . This corresponds to a Taylor expansion of the surface tension in the curved interface around the planar one. The lowest order corresponds to the planar surface tension, Eq. (191). The correction term, which is linear in the principal curvature $\kappa$ of the surface, is the curvature correction, which reads Vinas1998 ; douchin00 :
@@
$x$ $\sigma_{s}$ $\partial\sigma_{s}/\partial x$ $\sigma_{c}$ $\partial\sigma_{c}/\partial x$
0.025 1.23×$\times 10^{-4}$ 0.01154 -1.0×$\times 10^{-4}$ 0.125813
0.0375 9.91×$\times 10^{-4}$ 0.130393 0.001521 0.187507
0.05 0.003427 0.259847 0.005366 0.436
0.0625 0.007467 0.38018 0.012058 0.552647
0.075 0.012954 0.505293 0.018796 0.550167
0.0875 0.02018 0.652033 0.026064 0.636813
0.1 0.029337 0.817933 0.034889 0.76712
0.1125 0.040726 1.007927 0.04528 0.89944
0.125 0.054632 1.22098 0.057456 1.05268
0.1375 0.071351 1.458293 0.071684 1.22674
0.15 0.09119 1.719793 0.088219 1.424907
0.1625 0.114446 2.0058 0.107373 1.637987
0.175 0.141423 2.313307 0.129232 1.86716
0.1875 0.172345 2.63714 0.15409 2.10664
0.2 0.207386 2.969667 0.181881 2.34342
0.2125 0.246579 3.300753 0.212569 2.554153
0.225 0.289845 3.617873 0.245551 2.717393
0.2375 0.336904 3.90494 0.280327 2.84238
0.25 0.387305 4.152653 0.316336 2.896993
0.2625 0.440518 4.351273 0.352492 2.888607
0.275 0.495859 4.494633 0.388322 2.826707
0.2875 0.552645 4.580827 0.422924 2.705407
0.3 0.610126 4.60614 0.455786 2.5453
0.3125 0.66753 4.567813 0.486393 2.3485
0.325 0.724018 4.462667 0.514279 2.11932
0.3375 0.778609 4.234893 0.538593 1.68658
0.35 0.829358 3.860047 0.555488 0.963347
0.3625 0.875016 3.470013 0.563078 0.369047
0.375 0.916267 3.131693 0.565657 0.046033
0.3875 0.953379 2.8083 0.564588 -0.207787
0.4 0.986508 2.49228 0.560663 -0.41674
0.4125 1.015699 2.17876 0.554294 -0.597733
0.425 1.040985 1.86714 0.545816 -0.755767
0.4375 1.062381 1.556307 0.535472 -0.896807
0.45 1.079893 1.245467 0.523452 -1.024547
0.4625 1.093517 0.934433 0.509903 -1.141693
0.475 1.103253 0.623213 0.494946 -1.250187
0.4875 1.109096 0.311627 0.478678 -1.351613
0.5 1.111043 -1.2×$\times 10^{-4}$ 0.46118 -1.4472
[…]
The data for the surface tension may be fitted by the analytical fits. For the planar surface tension, the fit is given in Eq. (9). For the curvature correction, the fit is
@@
We can also determine the neutron reference position of the neutron interface $z_{n}$ in a similar way to that used to get $z_{p}$. Once the both interface positions, $z_{n}$ and $z_{p}$, are known, we can define the neutron skin thickness as $s_{n}=z_{n}-z_{p}$. This surface layer contains mainly neutrons, which can be considered as adsorbed onto the proton surface located at $z_{p}$. The number of neutrons in this layer per unit surface is given by $N_{s}/S=(n_{n0}-n_{no})s_{n}$ Vinas1998 ; douchin00 , where $S$ is the surface area, $n_{n0}$ is the neutron density in the bulk of nuclear matter.
There are additional contributions to the adsorbed neutrons due to the curvature of the proton surface, which per unit length $L$ are given by $N_{c}/L=(n_{n0}-n_{no})(b_{n}^{2}-b_{p}^{2}+s_{n}^{2})/2$ Vinas1998 ; douchin00 , where $b_{n}$ and $b_{p}$ are the surface width of the neutron and proton densities defined as
[…]
Appendix E Force acting on a single slab with neutron superflow in magnetic field
The basic assumption of the model developed in Kobyakov2018 , is that the system is a toroidal volume containing ideal 1N phase. The slabs are oriented along the field lines of the toroidal component $\mathbf{B}$ of the stellar magnetic induction, as schematically shown in Fig. 11 (a).
Figure 11: (Color online). Panel (a): A schematic representation (not to the scale) of the magnetic-equatorial cross-section of the star, highlighting the 1N phase located between the outer crust (when mean baryon density $n_{b}$ is smaller than the neutron drip density) and the core (when $n_{b}>n_{0}\eta_{\rm uni}^{*}$) and threaded by the toroidal lines of the stellar magnetic induction $\mathbf{B}$. Thin solid rectangle denoted as $\gamma$ represents the cut piece of the toroidal volume shown in Fig. 8. Panel (b): Enlarged view of 2-dimensional projection of 3-dimensional piece of volume $\gamma$ “cut” from panel (a), where $D$ is the height of the column discussed in Eq. (33). The planar nature of the system seen in panel (b) and in Fig. 8 is due to the very large curvature radius of the structure as compared with the structure lattice period, while the periodic boundary conditions in the plane parallel to the slabs, which are used in our calculations are essential for existence of the neutron superflow. The parameter $L$ was introduced in Eq. (52).
Let us now “cut” a piece of 1N phase in the inner crust, with spatial dimensions $D\times L^{2}$, as schematically shows thin rectangle in Fig. 11 (a). This rectangle is enlarged in Fig. 11 (b) in order to clearly define the physical system under consideration. By cutting a piece with $2r_{N}\ll L\ll R_{*}$, where $R_{*}$ is the stellar radius, we convert the physical problem of the inner crust shown in Fig. 11 (a) into an idealized model where the slabs are strictly flat as shown in Fig. 11 (b). This situation corresponds to the model shown in Fig. 8 (a).
We assume periodic boundary conditions for the nucleon number currents in the plane parallel to the slab, so that the neutron and proton flows parallel to the slab surface are possible. If the condition $d_{L}>\xi$ is satisfied, the slabs in Fig. 8 (a) are independent, and therefore it is sufficient to consider only a single slab, as shown in Fig. 8 (c).
Following Kobyakov2018 , we assume that the temperature is low enough, so that the superfluid nucleon densities are equal to the bulk nucleon densities $n_{\alpha}$, $(\alpha=p,n)$. The superconducting and superfluid S-wave order parameters $\sqrt{n_{\alpha}}\exp{\mathrm{i}\phi_{\alpha}}$ consist of the superfluid density $n_{\alpha}$ and the superfluid phase $\phi_{\alpha}$. The fluid specific momenta are given by $\mathbf{p}_{\alpha}=\hbar\nabla\phi_{\alpha}$. The energy density of the system shown in Fig. 8 (c) is given by
$w_{\rm tot}^{\rm matt}=w^{\rm nuc}+w_{\rm kin}^{\rm nuc}+w_{\rm Coul}^{\rm ep}+w^{\rm e},$ (196)
Here, $\boldsymbol{\pi}_{p}=\mathbf{p}_{p}-e\mathbf{A}/c$ is the gauge-invariant proton momentum, $\mathbf{A}$ is the electromagnetic vector potential, $\mathbf{w}=(\boldsymbol{\pi}_{p}-\mathbf{p}_{n})/m$ is the momentum lag, with $m=m_{p}\approx m_{n}$ and $\mathbf{B}=\nabla\times\mathbf{A}$.
The equations of motion generated from the energy functional with density given in Eq. (196) have the form of the continuity equations and the Newtown’s second law for each of the nucleon species Kobyakov2018 :
$\partial_{t}n_{\alpha}+\nabla\cdot\mathbf{J}_{\alpha}=0,$ (198)
where $\mathbf{J}_{\alpha}$ are the number currents of each species of nucleons. Combining the momentum equations we arrive at the expression for the total force acting under the given conditions.
Using the geometry shown in Fig. 8 (c), the equilibrium vector potential is $\mathbf{A}_{0}=-\mathbf{\hat{y}}B_{0}z$ and the unperturbed superfluid gauge-invariant momentum lag $\mathbf{w}_{0}$ is given by
$\mathbf{w}_{0}=(e/mc)\mathbf{\hat{y}}B_{0}z,$ (200)
see equation (35) in Kobyakov2018 . The force is zero in equilibrium, as also can be seen from Eq. (28) with $\mathbf{w}=\mathbf{w}_{0}$.
Let us study perturbation of the force as a result of perturbation of the neutron superflow. Notice that there is no back action of this force upon the source of the field $\mathbf{H}$, because this force is merely an extra fluid pressure gradient, which upon integration of the total force in the entire volume of the superconducting body, gives zero, as expected from the total momentum conservation. The integration gives zero because of the axial symmetry of the system as reveals Fig. 11 (a). In the coordinates shown in Fig. 8 (c), the relevant force with $L=1$ cm is given by
where the stress tensor is
Here, $\rho_{np}^{*}=mn_{np}-mn_{p}\theta_{p}-mn_{n}\theta_{n}$ is the entrainment mass density (in the Skyrme model, $\theta_{p}=\theta_{n}=0$), $n_{np}$ is the entrainment number density and $P$ is the total physical pressure (the hydrodynamic pressure). We evaluate the entrainment number density from (see Eq. (49) in KobyakovEtAl2017 )
$n_{np}=\frac{2}{n_{0}}\left(1-\frac{m}{m_{0}^{*}}\right)n_{n}n_{p},$ (203)
where $n_{p}$ and $n_{n}$ are the nucleon number densities in uniform matter, $m_{0}^{*}$ is the nucleon so-called isovector effective mass, which is given by (see Eq. (3.22) in ChabanatEtal1997 )
$\frac{m}{m_{0}^{*}}=1+\frac{m}{4\hbar^{2}}n_{0}\left[t_{1}\left(x_{1}+2\right)+t_{2}\left(x_{2}+2\right)\right].$ (204)
${m_{0}^{*}}=\left.\tilde{m_{q}}\right|_{n_{q}=0}.$ (205)
$P$ is generally defined as the material contribution to $\Pi_{ik}$ associated with the isotropic tensor $\delta_{ik}$. The hydrodynamic pressure is given by Kobyakov2018
$P=p-\rho_{np}^{*}\mathbf{w}^{2}/2,$ (206)
where $p$ is the pressure without the contributions due to the matter flows as given by equation (32) in Kobyakov2018 . The hydrostatic pressure $p$ is irrelevant in order to calculate the additional pressure due to the peculiar force and thus it is taken constant across the slab. The crustal stress appears in the problem as soon as we impose condition that the slab is static due to external forces at the boundary $\gamma$ of the domain, see Fig. 11. In this case $p$ becomes spatially dependent to compensate the additional pressure, and this affects the resulting total momentum flux at the domain boundary thus loading the crust until the resulting strain reaches a critical value.
It remains to evaluate the total force as a resultant of the forces applied to each of the slabs in the lattice (see Fig. 8 (a)). From independence of the slabs, the total force acting on a 1 ${\rm cm}^{2}\times D$ column of lasagna is a sum of (integer part of) ${D}/{2r_{c}}$ equal terms, or
where $D$ is the height of the column measured along the normal direction of the slabs.

### v2

@@ C.2 Variational equations
where $y$ is either of the variables from the set $\{n,\,x,\,n_{no},\,r_{B},\,u^{\rm bub}\}$.
@@ C.3 Optimum size of bubble
The extremum
$\left.\frac{\partial(w_{\rm tot}^{\rm bub})}{\partial r_{B}}\right|_{n,x,n_{no},u^{\rm bub}}=0$ (139)
@@
Notice a difference in signs in Eqs. (83) and (141). The solution to Eq. (141) is
@@ C.4 Chemical equilibrium
The extremum
$\left.\frac{\partial(w_{\rm tot}^{\rm bub})}{\partial x}\right|_{n,n_{no},r_{B},u^{\rm bub}}=0$ (147)
@@ C.5 Continuity of neutron chemical potential across the bubble interface
The extremum
$\left.\frac{\partial(w_{\rm tot}^{\mathrm{bub}})}{\partial n_{ni}}\right|_{N_{n},n_{pi},r_{B},u^{\mathrm{bub}}}=0$ (151)
@@
The partial derivative reads
$\left.\frac{\partial}{\partial n_{ni}}\right|_{N_{n},\,n_{pi},\,r_{B},\,u^{\mathrm{bub}}}=$ (154)
$\left.\frac{\partial}{\partial n}\right|_{x,\,n_{no},\,r_{B},\,u^{\mathrm{bub}}}+\frac{\delta x}{\delta n}\left.\frac{\partial}{\partial x}\right|_{n,\,n_{no},\,r_{B},\,u^{\mathrm{bub}}}$
$+\frac{\delta n_{no}}{\delta n_{ni}}\left.\frac{\partial}{\partial n_{no}}\right|_{n,\,x,\,r_{B},\,u^{\mathrm{bub}}}.$
@@
$\left.\frac{\partial}{\partial n_{ni}}\right|_{N_{n},\,n_{pi},\,r_{B},\,u^{\mathrm{bub}}}=$ (155)
$\left.\frac{\partial}{\partial n}\right|_{x,\,n_{no},\,r_{B},\,u^{\mathrm{bub}}}-\frac{x}{n}\left.\frac{\partial}{\partial x}\right|_{n,\,n_{no},\,r_{B},\,u^{\mathrm{bub}}}$
$-\frac{1-u^{\mathrm{bub}}}{u^{\mathrm{bub}}}\left.\frac{\partial}{\partial n_{no}}\right|_{n,\,x,\,r_{B},\,u^{\mathrm{bub}}}.$
@@
$+\frac{1-x}{(1-u^{\mathrm{bub}})n}\left.\frac{\partial(w_{\rm p.surf}^{\rm bub}+w_{\rm curv}^{\rm bub})}{\partial x}\right|_{r_{B},u^{\mathrm{bub}}}$
$+\frac{1}{(1-u^{\mathrm{bub}})n}\left.\frac{\partial(w_{\rm C+L}^{\rm bub})}{\partial x}\right|_{n,r_{B},u}.$ (159)
@@ C.6 Continuity of pressure across the bubble interface
The extremum
$\left.\frac{\partial(w_{\rm tot}^{\mathrm{bub}})}{\partial u^{\mathrm{bub}}}\right|_{N_{n},N_{no},x,V_{c}}=0$ (160)
implies that the energy density is minimized by choosing the optimal size of the bubble while keeping fixed the cell size, the number of neutrons, both inside and outside the bubble, and the proton fraction. As a result, the pressure inside the bubble is equal to that outside the bubble, $P_{i}^{\mathrm{bub}}=P_{o}^{\mathrm{bub}}$.
@@
The partial derivative reads
@@
$\times(w_{\rm p.surf}^{\mathrm{bub}}+w_{\rm cur}^{\mathrm{bub}}+w_{\rm C+L}^{\mathrm{bub}}),$
@@
$\times\left[\left(\frac{2+d}{d}+\frac{2u^{\mathrm{bub}}}{1-u^{\mathrm{bub}}}\right)f_{d}(u^{\mathrm{bub}})+u^{\mathrm{bub}}\partial_{u^{\mathrm{bub}}}f_{d}(u^{\mathrm{bub}})\right].$
@@ Appendix D Nuclear surface tension with curvature corrections
$x$ $\sigma_{s}\,[{\rm MeVfm}^{-2}]$ $\partial\sigma_{s}/\partial x$ $\sigma_{c}\,[{\rm MeVfm}^{-1}]$ $\partial\sigma_{c}/\partial x$
@@
0.0500 0.003427 0.2598 0.005366 0.4360
0.0625 0.007467 0.3802 0.01206 0.5526
0.0750 0.01295 0.5053 0.01880 0.5502
0.0875 0.02018 0.6520 0.02606 0.6368
0.1000 0.02934 0.8179 0.03489 0.7670
0.1125 0.04073 1.0079 0.04528 0.8994
0.1250 0.05463 1.2210 0.05746 1.0527
0.1375 0.07135 1.4583 0.07168 1.2267
0.1500 0.09119 1.7198 0.08822 1.4249
0.1625 0.1144 2.0058 0.1074 1.6380
0.1750 0.1414 2.3133 0.1292 1.8672
0.1875 0.1723 2.6371 0.1541 2.1066
0.2000 0.2074 2.9697 0.1819 2.3434
0.2125 0.2466 3.3007 0.2126 2.5542
0.2250 0.2898 3.6179 0.2456 2.7174
0.2375 0.3369 3.9049 0.2803 2.8424
0.2500 0.3873 4.1527 0.3163 2.8970
0.2625 0.4405 4.3513 0.3525 2.8886
0.2750 0.4959 4.4946 0.3883 2.8267
0.2875 0.5526 4.5808 0.4229 2.7054
0.3000 0.6101 4.6061 0.4558 2.5453
0.3125 0.6675 4.5678 0.4864 2.3485
0.3250 0.7240 4.4627 0.5143 2.1193
0.3375 0.7786 4.2349 0.5386 1.6866
0.3500 0.8294 3.8600 0.5555 0.9633
0.3625 0.8750 3.4700 0.5631 0.3690
0.3750 0.9163 3.1317 0.5657 0.04603
0.3875 0.9534 2.8083 0.5646 -0.2078
0.4000 0.9865 2.4923 0.5607 -0.4167
0.4125 1.0157 2.1788 0.5543 -0.5977
0.4250 1.0410 1.8671 0.5458 -0.7558
0.4375 1.0624 1.5563 0.5355 -0.8968
0.4500 1.0799 1.2455 0.5235 -1.0245
0.4625 1.0935 0.9344 0.5099 -1.1417
0.4750 1.1033 0.6232 0.4949 -1.2502
0.4875 1.1091 0.3116 0.4787 -1.3516
[…]
To obtain self-consistently the nuclear surface tension with curvature corrections, let us start considering a neutron-rich nucleus in equilibrium with a liquid of dripped neutrons. The problem of finding the density profile is formally similar to the one found in hot nucleus in equilibrium with the pure liquid phase. It is known that in this scenario the variational equations at given temperature and chemical potential possess two different solutions in both quantum bonche85 and semiclassical descriptions suraud87 . One of these solutions represents the nucleus-plus-liquid system and the another one to the pure liquid. Therefore, it seems natural to define the extensive quantities in the nucleus as the differences of their values in the nucleus-plus-liquid and the pure-liquid system.
To simplify this calculation, we simulate the nucleus as a half-space filled with uniform nuclear matter ($\tilde{z}<0$), while the other half-space is filled with the pure neutron liquid ($\tilde{z}>0$). As a consequence, in this geometry the energy per unit area, i.e. the surface tension, of nucleus constrained by the conservation of the neutron and protons numbers reads Vinas1998 :
[…]
The neutron and proton density profiles as well as the density of dripped neutrons are the solution of the Euler-Lagrange equations derived from Eq. (175):
where $\mu_{n}$ and $\mu_{p}$ are the neutron and proton chemical potentials computed at $\tilde{z}\to-\infty$, i.e. corresponding to the bulk phase in equilibrium with the dripped neutron liquid, which due to the equilibrium has the same neutron chemical potential $\mu_{n}$, and $\varepsilon_{no}=\varepsilon(n=n_{no},x=0)$.
The Euler-Lagrange equations (176) is a set of coupled second order differential equations, which are solved using the imaginary timestep method and provide self-consistent neutron and proton densities in LG and G phases. It is worth noting here that the surface tension defined by Eq. (175) corresponds to the surface contribution to the thermodynamical potential $\Omega=E-\mu_{n}N-\mu_{p}Z=-PV$ in LG phase, which is minimized in thermodynamic equilibrium between LG and G phases douchin00 .
The surface tension in Eq. (175) has been obtained under the assumption of a planar interface. In order to take into account that the interface is curved, we add an additional curvature correction that takes into account the curvature of the interface douchin00 . This corresponds to a Taylor expansion of the surface tension in the curved interface around the planar one. The lowest order corresponds to the planar surface tension, Eq. (175). The correction term, which is linear in the principal curvature of the surface, is the curvature correction, which reads Vinas1998 ; douchin00 :
$\sigma_{c}=\int^{+\infty}_{-\infty}d\tilde{z}[\varepsilon(\tilde{z})-\varepsilon_{no}-\mu_{n}(n_{n}(\tilde{z})-n_{no})-\mu_{p}n_{p}(\tilde{z})](\tilde{z}-z_{p}),$ (177)
where $z_{p}$ is the proton surface location, which is defined as $\int^{+\infty}_{-\infty}d\tilde{z}(\tilde{z}-z_{p})n^{\prime}_{p}(\tilde{z})=0$.
@@
The data for the surface tension may be fitted by analytical functions. For the planar surface tension, the fit is given in Eq. (9). For the curvature correction, the fit is